OpenStudy (amistre64):

Find the minimum sample size needed to estimate the population proportion: Margin of error: .015 Confidence level: 96% p ˆ ,q ˆ : unknown

OpenStudy (amistre64):

\[\hat p\ \text{and}\ \hat q\ \text{are unknown}\]

OpenStudy (amistre64):

i keep getting 4687, but it isnt one of my options .... either im doing it wrong, or the study guide is off

OpenStudy (amistre64):

\[n=\cfrac{(z_{\alpha/2})^2(.25)}{Error^2}\]

OpenStudy (blacksteel):

It's been a while since I did stats, but wouldn't that be \[\alpha = 0.96 => \alpha/2 = 0.48\]\[z_{0.48} = 0.6844\]\[n = 0.6844^2 * 025 / 0.015^2 \approx521\]

OpenStudy (blacksteel):

*0.25

OpenStudy (amistre64):

looks good; you zscore number is off tho ... I get it by doing (.5 + (.CI)/2)

OpenStudy (amistre64):

either way it gets to .48 with a zscore = 2.05

OpenStudy (blacksteel):

Yeah, that looks right - like I said, it's been a while. So then your equation should be\[n = (.9798)^2*0.25/0.015^2 \approx 1067\] right?

OpenStudy (amistre64):

my choices, albeit wrong as they may be, are : 4670 4519 3667 6669

OpenStudy (amistre64):

(.9798)^2 is where you are off at; the rest is good

OpenStudy (amistre64):

should be (2.05)^2 according to the text I have

OpenStudy (blacksteel):

Yeah, never mind, I'm an idiot. You're using the zscore, not the z table entry. When I run the numbers with (2.05)^2, I get 4670.

OpenStudy (amistre64):

4670 is good; i was using the calculators inverseNorm function which has a higher degree of accuracy than 2.05 which thew the answer off from the choices

OpenStudy (blacksteel):

Yup, that would do it.

OpenStudy (amistre64):

2.053748911 doesnt fit good in the answer choices :)